A certain metal is 20% tin. How many kilograms of this metal must be mixed with 80 kg of a metal that is 70% tin to get a mental that is 50% tin?
Answer by ptaylor(1894)

Let x=number of kilograms of 20% tin alloy needed
Now we know that the amount of pure tin in the 20% alloy (0.20x) plus the amount of pure tin in the 70% alloy (0.70*80) has to equal the amount of pure tin in the final mixture (0.50(80+x)). So, our equation to solve is:
0.20x+0.70*80=0.50(80+x) get rid of parens and simplify
0.20x+56=40+0.50x subtract 0.20x and also 40 from each side
0.20x-0.20x+56-40=40-40+0.50x-0.20x collect like terms
16=0.30x divide each side by 0.30
x=53.33 kilograms -----------amount of 20% alloy needed
CK
53.33*0.20+0.70*80=0.50*133.33
10.67+56=66.67
66.67=66.67
Hope this helps---ptaylor